OpenStudy (anonymous):

The height of a Phil Mickelson drive launched at time t = 0 can be patterned by the quadratic function \[h(t) = 15t - \frac34t^{2} + 4\]., where h(t) represents the height of the ball as a function of time t. Find the time t when the golf ball returns to its original launch height.

OpenStudy (anonymous):

when the golf ball returns to its original launch height is when h = 0 agree?

OpenStudy (anonymous):

so therefore, solve for t when h(t) = 0. remember to take the positive root

mathslover (mathslover):

just put t = 0

OpenStudy (anonymous):

i think when h=4 because at time t=0 h is 4

OpenStudy (foolaroundmath):

original height = h(0) = 4 So, what we need basically is h(t) = h(0) 15t - (3/4)t^2 +4 = 4 => (3/4)t^2 - 15t = 0 => t((3/4)t - 15) = 0 => t = 0 or t = 15/(3/4) = 20 t=0 is not true as it is the starting time Thus t = 20

OpenStudy (anonymous):

Oh ok. So you use 0 for h(t) and then just solve the quadratic?